The sum of the measures of two exterior angles of a quadrilateral is 300 and the other two exterior angles are congruent. Find the measure of each
Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


a,b,c,d = the four exterior angles

a+b = 300
c = d

For any polygon, the exterior angles add to 360 degrees
a+b+c+d = 360
(a+b)+c+c = 360
300+2c = 360
2c = 360-300
2c = 60
c = 60/2
c = 30
Each of the other two congruent exterior angles is 30 degrees.

The first two exterior angles (a and b) are unknown because we don't have enough information based solely on a+b = 300

Answer by ikleyn(52795)

.
The sum of the measures of two exterior angles of a quadrilateral is 300 and the other two exterior angles are congruent. Find the measure of each
~~~~~~~~~~~~~~~~~~~~~~


The problem formulation is incorrect and inaccurate from both
the mathematical and the regular human common sense points of view.

The correct formulation must sound this way

    The sum of the measures of two exterior angles of a quadrilateral is 300 
    and the other two exterior angles are congruent. Find the measure of each
    of the two the other exterior angles.

My correction is underlined in the third/last line.